# bio2Byte Tools deployment as a Python package and Galaxy tool to predict protein biophysical properties

### v1

@@ Abstract
Availability and Implementation: The suite is available on the Python Package Index (PyPI): https://pypi.org/project/b2bTools/ and Bioconda: https://bioconda.github.io/recipes/b2btools/README.html for Linux and macOS systems, with Docker images hosted on Biocontainers: https://quay.io/repository/biocontainers/b2btools?tab=tags&tag=latest and Docker Hub: https://hub.docker.com/u/bio2byte. Online deployments are available on Galaxy Europe: https://usegalaxy.eu/root?tool_id=b2btools_single_sequence and our online server: https://bio2byte.be/b2btools/. The source code can be found at https://bitbucket.org/bio2byte/b2btools_releases
@@ 2 Tools harmonisation and packaging
The tools included in this version of bio2Byte Tools were primarily sequence-based predictors, namely DynaMine [7], DisoMine [10], EfoldMine [9], AgMata [16] and PSPer [17], but also includes ShiftCrypt [18] for residue-level interpretation of NMR Chemical Shift data for proteins. Here, we focus on the description and integration of the sequence-based predictors, which facilitates their deployment in a single environment.
@@ 2.2 Programmatic usage and execution
To facilitate the execution of the predictors inside the package, their inner classes and methods were abstracted in a public wrapper class, named “SingleSeq” and illustrated in code snipped 1, which is the entry point for the users and acts as the interface between them and our predictors. This wrapper allows an easy parse of sequences in FASTA format, execution of the desired tools, storage for the outputs as a Python object dictionary, and export outputs and metadata as a JSON, CSV or TSV file (exemplified in code snippet 1 for Python usage and code snippet 2 for bash usage). An equivalent wrapper is also available to execute the tools on a multiple sequence alignment (MSA), named “MultipleSeq” and described in README.md, which maintains the aligned position of the amino acids and returns the bio2Byte Tools outputs indexed according to this alignment.
@@ 2.3 Testing and publishing
Our tools were then contained in a Python package. With every new version, this package is subject to exhaustive unit tests using the library pytest [21] to ensure codebase integrity and predictions consistency, i.e. we assert that the code modification does not alter the predictor’s outputs. We have automated the building and testing process by including a CD/CI pipeline in the package repository on Bitbucket. We run in parallel all the unit tests for every compatible Python version. Then the pipeline builds the Python package and tests its command-line execution with a simple FASTA input file.
If all tests are successful, a new version is manually published on PyPI and Bioconda, which additionally tests that all modules and sub-modules can be correctly imported and called. Once these tests are passed, the new version of bio2Byte Tools get incorporated in Bioconda’s index. As soon as this new version is included in Bioconda’s index, a new Docker image is automatically built and published on the Biocontainers containers registry. This image on Biocontainers is created only for Python 3.9, so we also create a set of Docker images for each new version in three different flavours for each Python from 3.7 to 3.12 (supplementary table 2). Each of these containers are subject to an additional iteration of the previously described unit tests upon creation. This ensures system-independent execution and reproducibility of results for every version of our software.
@@ 4 Future perspectives
Additionally, though this suite is able to use evolutionary information from an MSA to estimate the biophysical constraints of residue positions in a protein family, it does not provided additional interpretation of such results. We are currently working on an extra layer of interpretation for such information, for example to determine how well a protein fits within those biophysical constraints.
@@ 8 Funding
his work has been supported by the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement [813239 to J.G.-G. and A.D.]; Research Foundation Flanders (FWO) International Research Infrastructure [I000323N to W.V. and G.0328.16N to A.D.]; COST Action ML4NGP, CA21160, supported by COST (European Cooperation in Science and Technology).

### v2

@@ Abstract
Availability and Implementation: The suite is available on the Python Package Index (PyPI): https://pypi.org/project/b2bTools/ and Bioconda: https://bioconda.github.io/recipes/b2btools/README.html for Linux and macOS systems, with Docker images hosted on Biocontainers: https://quay.io/repository/biocontainers/b2btools?tab=tags&tag=latest and Docker Hub: https://hub.docker.com/u/bio2byte. Online deployments are available on Galaxy Europe: https://usegalaxy.eu/root?tool_id=b2btools_single_sequence and our online server: https://bio2byte.be/b2btools/. The source code can be found at https://bitbucket.org/bio2byte/b2btools_releases.
@@ 2 Tools harmonisation and packaging
The tools included in this version of bio2Byte Tools were primarily sequence-based predictors, namely DynaMine [7], DisoMine [10], EfoldMine [9], AgMata [16] and PSPer [17], which can estimate an array of metrics from a simple FASTA file, but also includes ShiftCrypt [18] for residue-level interpretation of NMR Chemical Shift data for proteins (Supplementary table 1). Here, we focus on the description and integration of the sequence-based predictors, which facilitates their deployment in a single environment. The steps described in this work result in further FAIRification of our tools [19], now robustly available for the scientific community.
@@ 2.2 Programmatic usage and execution
To facilitate the execution of the predictors inside the package, their inner classes and methods were abstracted in a public wrapper class, named “SingleSeq” and illustrated in code snippet 1, which is the entry point for the users and acts as the interface between them and our predictors. This wrapper allows an easy parse of sequences in FASTA format, execution of the desired tools, storage for the outputs as a Python object dictionary, and export outputs and metadata as a JSON, CSV or TSV file (exemplified in code snippet 1 for Python usage and code snippet 2 for bash usage). An equivalent wrapper is also available to execute the tools on a multiple sequence alignment (MSA), named “MultipleSeq” and described in README.md, which maintains the aligned position of the amino acids and returns the bio2Byte Tools outputs indexed according to this alignment.
@@ 2.3 Testing and publishing
Our tools were then contained in a Python package. With every new version, this package is subject to exhaustive unit tests using the library pytest [22] to ensure codebase integrity and predictions consistency, i.e. we assert that the code modification does not alter the predictor’s outputs. We have automated the building and testing process by including a CD/CI pipeline in the package repository on Bitbucket. We run in parallel all the unit tests for every compatible Python version. Then the pipeline builds the Python package and tests its command-line execution with a simple FASTA input file (supplementary figure 1).
If all tests are successful, a new version is manually published on PyPI and Bioconda, which additionally tests that all modules and sub-modules can be correctly imported and called. Once these tests are passed, the new version of bio2Byte Tools get incorporated in Bioconda’s index. As soon as this new version is included in Bioconda’s index, a new Docker image is automatically built and published on the Biocontainers containers registry. This image on Biocontainers is created only for Python 3.10, so we also create a set of Docker images for each new version in three different flavours for each Python from 3.7 to 3.12 (supplementary table 3). Each of these containers are subject to an additional iteration of the previously described unit tests upon creation. This ensures system-independent execution and reproducibility of results for every version of our software.
@@ 4 Future perspectives
Additionally, though this suite is able to use evolutionary information from an MSA to estimate the biophysical constraints of residue positions in a protein family, it does not provide additional interpretation of such results. We are currently working on an extra layer of interpretation for such information, for example to determine how well a protein fits within those biophysical constraints.
@@ 8 Funding
This work has been supported by the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement [813239 to J.G.-G. and A.D.]; Research Foundation Flanders (FWO) International Research Infrastructure [I000323N to W.V. and G.0328.16N to A.D.]; COST Action ML4NGP, CA21160, supported by COST (European Cooperation in Science and Technology).
@@
Table 1: Tools description. Simple description of every tool contained in bio2Byte tools.
Predictor Usage
DynaMine Fast predictor of protein backbone dynamics using only sequence information as input. The current version also predicts side-chain dynamics and secondary structure predictors using the same principle.
DisoMine Predicts protein disorder with recurrent neural networks not directly from the amino acid sequence, but instead from more generic predictions of key biophysical properties, here protein dynamics, secondary structure and early folding.
EfoldMine Predicts from the primary amino acid sequence of a protein, which amino acids are likely involved in early folding events.
AgMata Single-sequence based predictor of protein regions that are likely to cause beta-aggregation.
PSPer PSP (Phase Separating Protein) predicts whether a protein is likely to phase-separate with a particular mechanism involving RNA interacts (FUS-like proteins). It will highlight protein regions that are involved mechanistically, and provide an overall score.
ShiftCrypt Auto-encoding NMR chemical shifts from their native vector space to a residue-level biophysical index.
@@
Figure 1: Pytest process diagram. The package testing process starts from a single Git repository, which then creates an environment per Python version from 3.7 to 3.12. For each environment, a series of tests on the source code are executed, namely the execution of single sequence, multiple sequence, parsers, and ShiftCrypt/NMR modules. Then the outputs get compared to ground results values and assessed for differences beyond float point errors (sum of absolute differences across all predicted metrics and residues, proteins and residues smaller than 0.001). A python package is then created and single sequence and multiple sequence modules are newly tested. If all tests across all Python versions are satisfactory, the python package will be uploaded to PyPI and to Bioconda. Then, an array of Docker images are created and uploaded to Docker Hub, following the settings in supplementary table 3. Finally, Bioconda automatically deposits a docker container with the Python 3.10 version of the package in Biocontainers, which can then be deployed into Galaxy.